OpenStudy (anonymous):

Please help me!!!! Verify each trigonometric equation by substituting identities to match the right hand side of the equation to the left hand side of the equation. (tan(X)+tan(y))/1-tan(x)tan(y)=cot(x)+cot(y)/cot(x)cot(y)-1

OpenStudy (campbell_st):

the left side of the equation is \[\tan(x + y)\] the right hand side is \[\frac{1}{\cot(x + y)} = \frac{\cot(x)+\cot(y)}{\cos(x)\cot(y) -1}\] so 1/cot = tan therefore left hand side = right hand side